find rate of change of f(x)=-2x^2+5 from 2 to 4 from 1to2 from -3to 0
Answer by Alan3354(69443)

find rate of change of f(x)=-2x^2+5 from 2 to 4 from 1to2 from -3to 0
---------
f(x)=-2x^2+5
f(4) = -2*4^2 + 5 = -32+5 = -27
f(2) = -2*2^2 + 5 = -8 +5 = -3
ROC = (-27 -(-3))/(4-2) = -24/2 = -12
----------
You might notice that the 5 has no effect, and could be left out of the calculations.
====================
Do the other 2 the same way.